OpenStudy (anonymous):

Find the zeros of f(x) = 16x^4 - 81?

OpenStudy (gorv):

16x^4 - 81=0 (2x)^4=3^4 2x=3 x=3/2

OpenStudy (gorv):

x=+3/2 and -3/2

OpenStudy (gorv):

@jjosephs23